OpenStudy (toxicsugar22):

The center of a circle is the point C(2,4), and P(-1,6) is a point on the circle. Find the equation of the line tangent to the circle at the point P.

OpenStudy (anonymous):

use distanace formula to find radius

OpenStudy (toxicsugar22):

Cnu help me

OpenStudy (anonymous):

D from C to P = \[\sqrt{3^2+2^2} = \sqrt{13}\] \[(x-2)^2 + (y-4)^2 = 13\]

OpenStudy (toxicsugar22):

ok is this the answer

OpenStudy (anonymous):

Yes :D

OpenStudy (toxicsugar22):

i think we write in in slope intecept form

OpenStudy (toxicsugar22):

so what would that be

OpenStudy (anonymous):

You can only write a line in slope-int form, not a circle :D

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

thanks can u help me with this one

OpenStudy (toxicsugar22):

sketch the triangle with vertics A(-5,12), B(4,-2) and C(1,-6) Find the slope of the side AC. Find the slope of the altitude form point B to side AC fIND an equation for the line that includes the altitude from point B to side AC.